OpenStudy (anonymous):

solution of this equation and how to do it? -2y+5+5y=14

OpenStudy (mertsj):

Combine like terms on the left first.

OpenStudy (anonymous):

-7y

OpenStudy (mertsj):

-2y+5y=3y

OpenStudy (anonymous):

why 3y?

OpenStudy (mertsj):

Because if you have 5 dollars, and spend 2 dollars, you have 3 dollars left.

OpenStudy (anonymous):

Hi :) let meh explain this equation to you cudder

OpenStudy (mertsj):

So 5d+(-2d)=3d and -2d + 5d = 3d and -2y+5y=3y

OpenStudy (anonymous):

so is this when you use inverse operation?

OpenStudy (mertsj):

Not to combine the like terms.

OpenStudy (anonymous):

ok give it a try

OpenStudy (mertsj):

Now you have 3y+5=14 So now use the inverse operation of adding 5 which is subtract 5 from both sides of the equation.

OpenStudy (anonymous):

ok

OpenStudy (mertsj):